# The Optimal Paper Moebius Band

### 28 Dec 2023 v9

@@ 1 Introduction
  • For immersed paper Moebius bands one has $\lambda>\pi/2$. In particular, this bound holds for embedded paper Moebius bands. Moreover, one can find a sequence of immersed examples with the aspect ratio converging to $\pi/2$. These examples are not embedded.
  • One can find a sequence of embedded paper Moebius bands with the aspect ratio converging to $\sqrt{3}$. These examples converge to the triangular Moebius band. The triangular Moebius band itself does not count as an embedded smooth paper Moebius band.
The last line of [HW] states the conjecture that $\lambda>\sqrt{3}$ for an embedded paper Moebius band.
@@
Let $I_{n}:M_{\lambda_{n}}\to\Omega_{n}$ be a sequence of smooth embedded paper Moebius bands such that $\lambda_{n}\to\sqrt{3}$. Then, up to isometry, $I_{n}$ converges uniformly to the map giving the triangular Moebius band.
Now let me explain the strategy of the proof. Let $\Omega$ be a smooth embedded paper Moebius band. A bend on $\Omega$ is a line segment which cuts across $\Omega$ and has its endpoints in the boundary. We say that a $T$-pattern on $\Omega$ is a pair of bends which lie in perpendicular intersecting lines. We call the $T$-pattern embedded if the two bends are disjoint. In §2.1 we prove
@@
Our proof is essentially topological. It is well known that $\Omega$ has a partition into continuously varying bends. We apply a variant of the Borsuk-Ulam Theorem to a pair of functions describing the geometry of pairs of bends in this partition; the simultaneous vanishing of these functions gives us our embedded $T$-pattern. Our proof, slightly modified, would also show that an immersed paper Moebius band has a $T$-pattern.
@@
The basic idea of the proof is to cut $\Omega$ open along the bends comprising the $T$ pattern and then to explore the geometry of the situation.
The Main Theorem is an immediate consequence of Lemma T and Lemma G. The proof of the Triangular Limit Theorem, which we give in §2.3, amounts to examining what our proof of Lemma G says about a minimizing sequence.
The proofs of our theorems are done after §2, but I also include some more material. In §3, which is not needed for our proofs, I will enlarge on various topics related to the results proved here. In §4, an appendix, I give a self-contained and elementary proof of the result that a smooth embedded paper Moebius band has a continuous partition into bends. I found the proofs in the literature not so easy to understand.
Here is some additional context for this work. The work here supersedes my earlier paper [S1] and also is independent from it, but nonetheless it is an outgrowth of [S1]. In [S1] I improved the lower bound $\lambda>\pi/2$ in the embedded case to a bound $\lambda>\lambda_{1}$ for some complicated number $\lambda_{1}\in(\pi/2,\sqrt{3})$. I did this by first proving Lemma T (with some side hypotheses) and then deducing that $\lambda\geq\phi$, the golden ratio. This amounts to taking $t=0$ in Equation 4 below. Then I solve an optimization problem to get a lower bound $\lambda_{1}\in(\phi,\sqrt{3})$.
Embarrassingly, I discovered recently that I made an error in setting up the optimization problem in [S1]. I mistakenly assumed that when you cut $M_{\lambda}$ open along an embedded line segment which joins two points in $\partial M_{\lambda}$ you get a parallelogram rather than a trapezoid. This idiotic mistake caused me to miscalculate $\lambda_{1}$. So, all I can conclude from [S1] is that $\lambda_{1}\geq\phi$. However, I was amazed and delighted to discover that when I did the optimization problem correctly I got $\lambda_{1}=\sqrt{3}$ right on the nose! This corrected calculation is Lemma G above.
The topic of paper Moebius bands is adjacent to a number of different subjects. The paper [GKS] considers the related question of tying a piece of rope into a knot using as little rope as possible. See [DDS] for further results. One could view these rope knot questions as variants of the Halpern-Weaver Conjecture in a different category. Indeed, our Lemma T seems quite related in spirit to the quadrisecent idea in [DDS]. I will say a bit more about this in §3.2.
@@
Here is one more thing I’d like to mention. Some readers might find this paper hard to read because I do not include much background information. I have subsequently written a longer and friendlier account [S2], aimed at university students and perhaps advanced high school students. This paper is available on my Brown University website. My website also has some informal notes which are even more elementary.
@@ 2.1 Proof of Lemma T
Proof: Since $g(x,y)=0$ the vectors $\overrightarrow{u}$ and $\overrightarrow{v}$ are orthogonal. Hence $\overrightarrow{n}=\overrightarrow{u}\times\overrightarrow{v}$ is nonzero. By construction $u$ and $v$ and the segment $\overline{m_{u}m_{v}}$ all lie in planes orthogonal to $\overrightarrow{n}$. But then they all lie in the same plane orthogonal to $\overrightarrow{n}$. In short, $u$ and $v$ are co-planar. The bends are disjoint because they are in the same partition. $\spadesuit$
To prove Lemma T, we just have to prove that $g$ and $h$ simultaneously vanish somewhere in $\Upsilon$. Suppose not. Since $|g|\geq 1$ on $\partial\overline{\Upsilon}$, we can say that $g$ and $h$ do not simultaneously vanish on $\overline{\Upsilon}$. Let $S^{1}$ be the unit circle. Let $A=(f,g)$ and $B=A/\|A\|$. Then $B:\overline{\Upsilon}\to S^{1}$ is well-defined and continuous. $B$ maps one component of $\partial\overline{\Upsilon}$ to $(1,0)$ and the other to $(-1,0)$.
Consider any path $\gamma$ which connects a point in one component of $\partial\overline{\Upsilon}$ to a point in the other. The image $B(\gamma)$, always oriented from $(1,0)$ to $(-1,0)$, winds some half integer $w(\gamma)$ times around the origin. All choices of $\gamma$ are homotopic to each other relative to $\partial\overline{\Upsilon}$. Thus $w(\gamma)$ is independent of $\gamma$. In particular, $w(\Sigma(\gamma))=w(\gamma)$. However, $B\circ\Sigma=-B$. So, as Figure 2 illustrates, when we orient $B(\Sigma(\gamma))=-B(\gamma)$ from $(1,0)$ to $(-1,0)$, the winding number is $-w(\gamma)$. This contradiction completes the proof.
@@ 3.1 Lemma G
    There exist $2$ segments $v,w$ in the partition such that $I(v)$ and $I(w)$ are disjoint and lie in perpendicular intersecting lines.
@@ 3.2 Lemma T
Borsuk-Ulam Theorem: The proof I give of Lemma T is quite reminiscent of the proof of the Borsuk-Ulam Theorem. Indeed, Jeremy Kahn pointed out to me that the endgame of my proof really is the Borsuk-Ulam proof in disguise. To see this, note that we obtain the $2$-sphere $S^{2}$ by crushing each component of $\partial\overline{\Upsilon}$ to a point. Then $B$ induces a map $S^{2}\to S^{1}$ with $B\circ\Sigma=-\Sigma$. The map $\Sigma$, which is a glide reflection on $\overline{\Upsilon}$, acts on $S^{2}$ as the antipodal map.
In this context, it is more natural to redefine the vectors $\overrightarrow{u}$ and $\overrightarrow{v}$ to be the unit vectors parallel to the orientations of $u$ and $v$. Once this is done, the functions $g$ and $h$ themselves restrict to our quotient $S^{2}$ and we really have the exact conditions for the Borsuk-Ulam Theorem.
@@ 3.4 More Twists
One can define an $n$-twisted paper strip in the obvious way. When $n$ is odd, these are paper Moebius bands and when $n$ is even these are paper cylinders. Let $\lambda_{n}$ be the infimal value of aspect ratios of $n$-twist embedded paper strips. Our Main Theorem combines with the result in [HW] to say that $\lambda_{1}=\sqrt{3}$. The results in [S3] say that $\lambda_{2}=2$. We conjecture in [BS] that $\lambda_{3}=3$. Noah Montgomery has a construction showing that $\lambda_{n}$ grows at most like $\sqrt{n}$. Recently, Aidan Hennessey showed me a construction which seems to prove that $\lambda_{n}<8$ for all $n$. I hope that Aidan presents this construction in a forthcoming paper.
@@ 4.1 The Proof Modulo a Detail
On the bottom of p. 46 of [HW], Halpern and Weaver say that the result of Lemma 4.2 is well known. They cite the references [CL], [HN], and [St]. More precisely, Lemma 4.2 is a special case of the two essentially identical results, [CL, p. 314, Lemma 2] and [HN, §3, Lemma 2]. These results and proofs are done in a general multi-dimensional setting. Below I give an elementary and geometric proof tailored to the $2$-dimensional case.
It follows immediately from Lemma 4.2 that $U$ has a continuous partition into bends. The uniqueness implies the continuity. Let $\tau$ be a component of $\Omega-U$. If $\tau$ has empty interior then $\tau$ is a line segment, the limit of a sequence of bends. In this case $\tau$ is also a bend. Suppose $\tau$ has non-empty interior. The Gauss map is constant on $\tau$ and hence $\tau$ lies in a single plane. Two sides of $\tau$, opposite sides, lie in $\partial\Omega$ and are straight line segments. The other two sides of $\tau$, the other opposite sides, are bends. Thus $\tau$ is a planar trapezoid. But then we can extend our bend partition across $\tau$ by simply choosing any continuous family of segments on $\tau$ that interpolates between the two bends in its boundary. Doing this construction on all such components, we get our continuous partition of $\Omega$ into bends.
@@
Now we turn to the proof of Lemma 4.2. Let $U\subset\Omega^{o}$ as above. Let $p\to n_{p}$ be a local choice of the Gauss map. We can rotate and translate so that near the origin $U$ is the graph of a function
@@
Here $C>0$ is some constant. The normal vector at the origin is $n_{0}=(0,0,1)$. The vector $v_{0}=(1,0,0)$ lies in the kernel of $dn_{0}$. Let $w_{0}=v_{0}\times n_{0}=(0,1,0)$. Let $\Pi_{0}$ be the plane spanned by $w_{0}$ and $n_{0}$. The image of $\Pi_{0}\cap U$ under the Gauss map is (near $n_{0}$) a smooth regular curve tangent to $w_{0}$ at $n_{0}$.
@@
$v\to n_{p},\hskip 30.0ptp\to v_{p},\hskip 30.0ptv\to w_{p}=v_{p}\times n_{p}.$ (8)

### 31 Dec 2023 v10

@@ 1 Introduction
  • For smooth immersed paper Moebius bands one has $\lambda>\pi/2$. Moreover, for any $\epsilon>0$ one can find an immersed example with $\lambda=\pi/2+\epsilon$.
  • There exists some $\epsilon_{0}>0$ such that $\lambda>\pi/2+\epsilon_{0}$ for a smooth embedded paper Moebius band. This $\epsilon_{0}$ is not an explicit constant.
Halpern and Weaver give examples of smooth embedded paper Moebius bands with $\lambda=\sqrt{3}+\epsilon$ for any $\epsilon>0$. As they point out, similar examples appear in [Sa]. These examples converge to the triangular Moebius band from Figure 1 as $\epsilon\to 0$. The last line of [HW] states the conjecture that $\lambda>\sqrt{3}$ for an embedded paper Moebius band.
@@
Let $I_{n}:M_{\lambda_{n}}\to\Omega_{n}$ be a sequence of smooth embedded paper Moebius bands such that $\lambda_{n}\to\sqrt{3}$. Then, up to isometry, $I_{n}$ converges uniformly to the triangular Moebius band map.
Let me explain the strategy of the proofs. Let $\Omega$ be a smooth embedded paper Moebius band. A bend on $\Omega$ is a line segment which cuts across $\Omega$ and has its endpoints in the boundary. We say that a $T$-pattern on $\Omega$ is a pair of bends which lie in perpendicular intersecting lines. We call the $T$-pattern embedded if the two bends are disjoint. In §2.1 we prove
@@
Our proof is topological. It is well known that $\Omega$ has a partition into continuously varying bends. We apply (essentially) the Borsuk-Ulam Theorem to a pair of functions describing the geometry of pairs of bends in this partition; the simultaneous vanishing of these functions gives us our embedded $T$-pattern. Our proof, slightly modified, would also show that an immersed paper Moebius band has a $T$-pattern.
@@
The basic idea of the proof is to cut $\Omega$ open along one of the bends comprising the $T$ pattern and then to solve an optimization problem which involves mapping a trapezoid into space with certain constraints.
The Main Theorem is an immediate consequence of Lemma T and Lemma G. The proof of the Triangular Limit Theorem, given in §2.3, amounts to examining what our proof of Lemma G says about a minimizing sequence.
The proofs are done after §2, but I also include some more material. In §3 I will comment on the proofs and discuss some related topics. In §4, an appendix, I give a self-contained and elementary proof of the result that a smooth embedded paper Moebius band has a continuous partition into bends. I found the proofs in the literature not so easy to understand.
The ideas in this paper are an outgrowth of my earlier paper [S1]. In [S1] I prove a version of Lemma T (with some side hypotheses) and then deduce from the $t=0$ case of Equation 4 below that $\lambda>\phi=(1+\sqrt{5})/2$ in the embedded case. I solved an optimization problem akin to Lemma G in [S1] in order to further improve the bound, but I made an idiotic mistake: I thought that when you cut $M_{\lambda}$ open along a bend you get a parallelogram rather than a trapezoid. This mistake invalidates the further bound, and all I can conclude from [S1] is that $\lambda>\phi$. I was amazed and delighted to discover that the correctly done optimization problem, namely Lemma G, gets the sharp bound. The work here supersedes [S1] and is completely independent from it.
Some readers might find this paper hard to read because I do not include much background information. I have subsequently written a longer and friendlier account [S2], aimed at university students and perhaps advanced high school students. This paper is available on my Brown University website. My website also has some informal notes which are even more elementary.
Here is some additional context. The topic of paper Moebius bands is adjacent to a number of different subjects. The paper [GKS] considers the related question of tying a piece of rope into a knot using as little rope as possible. See [DDS] for further results. One could view these rope knot questions as variants of the Halpern-Weaver Conjecture in a different category. Indeed, our Lemma T seems quite related in spirit to the quadrisecent idea in [DDS]. I will say a bit more about this in §3.2.
@@ 2.1 Proof of Lemma T
Proof: Since $g(x,y)=0$ the vectors $\overrightarrow{u}$ and $\overrightarrow{v}$ are orthogonal. Hence $\overrightarrow{n}=\overrightarrow{u}\times\overrightarrow{v}$ is nonzero. By construction $u$ and $v$ and the segment $\overline{m_{u}m_{v}}$ all lie in planes orthogonal to $\overrightarrow{n}$. Since $u\cup v\cup\overline{m_{u}m_{v}}$ is connected, this union lies in the same plane orthogonal to $\overrightarrow{n}$. In short, $u$ and $v$ are co-planar. The bends are disjoint because they are in the same partition. $\spadesuit$
To prove Lemma T, we just have to prove that $g$ and $h$ simultaneously vanish somewhere in $\Upsilon$. Suppose not. Since $|g|\geq 1$ on $\partial\overline{\Upsilon}$, we can say that $g$ and $h$ do not simultaneously vanish on $\overline{\Upsilon}$. Let $S^{1}$ be the unit circle. Let $A=(g,h)$ and $B=A/\|A\|$. Then $B:\overline{\Upsilon}\to S^{1}$ is well-defined and continuous. $B$ maps one component of $\partial\overline{\Upsilon}$ to $(1,0)$ and the other to $(-1,0)$.
Consider a path $\gamma$ which connects a point in one component of $\partial\overline{\Upsilon}$ to a point in the other. The image $B(\gamma)$, when oriented from $(1,0)$ to $(-1,0)$, winds some half integer $w(\gamma)$ times around the origin. All choices of $\gamma$ are homotopic to each other relative to $\partial\overline{\Upsilon}$. Thus $w(\gamma)$ is independent of $\gamma$. In particular, $w(\Sigma(\gamma))=w(\gamma)$. However, $B\circ\Sigma=-B$. So, when we orient $B(\Sigma(\gamma))=-B(\gamma)$ from $(1,0)$ to $(-1,0)$, as illustrated in Figure 2, the winding number is $-w(\gamma)$. This contradiction completes the proof.
@@ 3.1 Lemma G
    There exist $2$ segments $v,w$ in the partition and a linear projection $\phi$ onto a plane such that $\phi\circ I$ is distance non-decreasing on each of $v$ and $w$ and such that $\phi\circ I(v)$ and $\phi\circ I(w)$ are disjoint and perpendicular.
@@ 3.2 Lemma T
Borsuk-Ulam Theorem: The proof I give of Lemma T is quite close to a proof of the Borsuk-Ulam Theorem. Indeed, Jeremy Kahn pointed out to me that the endgame of my proof really is the Borsuk-Ulam proof in disguise. To see this, note that we obtain the $2$-sphere $S^{2}$ by crushing each component of $\partial\overline{\Upsilon}$ to a point. Then $B$ induces a map $S^{2}\to S^{1}$ with $B\circ\Sigma=-B$. The map $\Sigma$, which is a glide reflection on $\overline{\Upsilon}$, acts on $S^{2}$ as the antipodal map.
There is a very nice way to do all this explicitly. We think of $\overline{\Upsilon}$ as the set of equivalence classes of points $(x,y)\in\mbox{\boldmath{$R$}}^{2}$ such that $x\leq y\leq x+1$. The equivalence relation is $(x,y)\sim(x+k,y+k)$ for all $k\in\mbox{\boldmath{$Z$}}$. Here is a continuous surjection from $\overline{\Upsilon}$ to $S^{2}$ which maps the boundary components to $(0,0,\pm 1)$ and is a homeomorphism from $\Upsilon$ to $S^{2}-\{(0,0,\pm 1)\}$:
$(x,y)\to(\sqrt{1-z^{2}}\cos(\theta),\sqrt{1-z^{2}}\sin(\theta),z),$
$\theta=\pi(x+y),\hskip 30.0ptz=2x-2y+1.$ (7)
The action of $\Sigma$ on this version of $\Upsilon$ maps the class of $(x,y)$ to the class of $(y,x+1)$. The map in Equation 7 respects the equivalence relation and conjugates $\Sigma$ to the antipodal map.
We now redefine the vectors $\overrightarrow{u}$ and $\overrightarrow{v}$ to be the unit vectors parallel to the orientations of $u$ and $v$. Once this is done, the functions $g$ and $h$ themselves extend to $S^{2}$ and satisfy $g(-P)=-g(P)$ and $h(-P)=-h(P)$ for all $P\in S^{2}$. We then conclude from the Borsuk-Ulam Theorem that there is some $P\in S^{2}$ such that $g(P)=h(P)=0$. Since $g((0,0,\pm 1))=\pm 1$ we conclude that $P\not=(0,0,\pm 1)$. Hence $P$ actually corresponds to a $T$-pattern.
@@ 3.4 More Twists
One can define an $n$-twisted paper strip in the obvious way. When $n$ is odd, these are paper Moebius bands and when $n$ is even these are paper cylinders. Let $\lambda_{n}$ be the infimal value of aspect ratios of $n$-twist embedded paper strips. Our Main Theorem combines with the result in [HW] to say that $\lambda_{1}=\sqrt{3}$. The results in [S3] say that $\lambda_{2}=2$. We conjecture in [BS] that $\lambda_{3}=3$. Noah Montgomery has a construction showing that $\lambda_{n}$ grows at most like $\sqrt{n}$. Recently, Aidan Hennessey showed me a construction which seems to prove that $\lambda_{n}<8$ for all $n$.
@@ 4.1 The Proof Modulo a Detail
Lemma 4.2 is a special case of the two essentially identical results, [CL, p. 314, Lemma 2] and [HN, §3, Lemma 2]. These results and proofs are done in a general multi-dimensional setting. Below I give an elementary and geometric proof tailored to the $2$-dimensional case.
It follows immediately from Lemma 4.2 that $U$ has a continuous partition into bends. The uniqueness implies the continuity. Let $\tau$ be a component of $\Omega-U$. If $\tau$ has empty interior then $\tau$ is a line segment, the limit of a sequence of bends. In this case $\tau$ is also a bend. Suppose $\tau$ has non-empty interior. The Gauss map is constant on $\tau$ and hence $\tau$ lies in a single plane. Two sides of $\tau$, opposite sides, lie in $\partial\Omega$ and are straight line segments. The other two sides of $\tau$, the other opposite sides, are bends. Thus $\tau$ is a planar trapezoid. But then we can extend our bend partition across $\tau$ by simply choosing any continuous family of segments on $\tau$ that interpolates between the two bends in its boundary. Indeed, there is a canonical extension: If the two bends of $\partial\tau$ lie on parallel lines we can take the interpolating bends to lie on parallel lines. If they lie on intersecting lines, we can take the interpolating bends to lie on lines which all go through the intersection point. Doing this construction on all such components, we get our continuous partition of $\Omega$ into bends.
@@
Let $U\subset\Omega^{o}$ as above. Let $p\to n_{p}$ be a local choice of the Gauss map. We can rotate and translate so that near the origin $U$ is the graph of a function
@@
Here $C>0$ is some constant. The normal vector at the origin is $n_{0}=(0,0,1)$. The vector $v_{0}=(1,0,0)$ lies in the kernel of $dn_{0}$. Let $w_{0}=(0,1,0)$. Let $\Pi_{0}$ be the plane spanned by $w_{0}$ and $n_{0}$. The image of $\Pi_{0}\cap U$ under the Gauss map is (near $n_{0}$) a smooth regular curve tangent to $w_{0}$ at $n_{0}$. The sign depends on the choice of local Gauss map.
@@
$p\to n_{p},\hskip 30.0ptp\to v_{p},\hskip 30.0ptp\to w_{p}=v_{p}\times n_{p}.$ (9)